# Runyang You1    Yongqi Li1   Xinyu Lin2   Xin Zhang1 Wenjie Wang3   Wenjie Li1   Liqiang Nie4 1

### v1

@@ Appendix B Baselines
Table 3: Dataset Statistics.
Dataset Users Items Density Interactions Train Val Test
Video Games 29,230 10,144 0.031% 63,502 50,801 6,350 6,351
CDs and Vinyl 7,701 12,024 0.023% 13,435 10,748 1,343 1,344
Musical Instruments 15,656 10,320 0.031% 34,373 27,498 3,437 3,438
Appendix C Implementation Details
@@
[…]
$\text{R}^{2}\text{ec}$ Optimization.
@@
Appendix D Overall Pipeline
@@
7:     Sample a training batch =$\mathcal{B}=\{(u,v^{+})\}\sim\mathcal{D}$
@@
Appendix E Prompt
User Prompt Analyze in depth and finally recommend next {category} I might purchase inside <answer> and </answer>  . For example, <answer> a product </answer>  . Below is my historical {category} purchases and ratings (out of 5): {for hist in purchase_histories %} {{hist.time_delta} ago: [{hist.item_title}] ({hist.rating}) %}
Item Prompt Summarize key attributes of the following {category} inside <answer> and </answer>  : {for key, attr in item.meta %} {{key}: {attr} %}
@@ Appendix F Limitations
Appendix G Analysis on Inference Efficiency
Table 4: Average per-request inference latency (mean) of our unified model $\text{R}^{2}\text{ec}$ versus the two-stage LangPtune pipeline, measured across varying batch sizes.
batch size $\text{R}^{2}\text{ec}$ LangPTune
1 16725.54 $\pm$ 2636.84 19030.95 $\pm$ 2753.81
4 3980.77 $\pm$ 77.13 4625.045 $\pm$ 67.00
8 2220.88 $\pm$ 32.00 2375.64 $\pm$ 69.476723
To quantify the efficiency gains of our unified model $\text{R}^{2}\text{ec}$ over the two-stage conventional pipeline, we benchmark its inference latency against LangPtune [15]. Inferences were performed on one NVIDIA RTX 3090 GPU using identical input prompts, maximum token lengths, and candidate sets. We executed 100 queries at batch sizes of 1, 4, and 8. For each batch, per-request latency was calculated by dividing the total duration by the batch size. The mean $\pm$ standard deviation for each configuration is reported in Table 4. $\text{R}^{2}\text{ec}$ achieves approximately 10% reduction in average per-request inference time, demonstrating the superior efficiency of our model architecture compared to the conventional two-stage approach. Notably, the efficiency gap narrows as batch size increases, since parallelization spreads computation costs over more requests. In addition, the low standard deviations demonstrate that $\text{R}^{2}\text{ec}$’s inference speed is reliably stable.
Appendix H Analysis on Case Study
To illustrate domain-specific and backbone-specific reasoning behaviors, we analyze representative cases from the CDs and Video Games datasets (Figures 5,6,7,8) using Gemma and Qwen as backbones. 1) On the CDs dataset, both models tend to reason by discussing several potentially interesting items and explicitly enumerating their features and genres, before providing a final recommendation. This behavior highlights the models’ ability to leverage domain knowledge acquired during pretraining and reinforced through RecPO, for better recommendation. 2) On the Video Games dataset, the models shift towards inferring user intent from the purchase sequence, speculating on the likely category or intended use of the next purchase. This demonstrates a shift from feature-based enumeration to intent-driven reasoning as the domain changes. 3) Notably, Gemma generally produces shorter and more concise reasoning sequences compared to Qwen, which often generates more elaborate justifications. These patterns indicate that, while both models benefit from explicit reasoning, their generation style and reliance on pretrained knowledge can manifest differently across domains and backbone choices.

### v3

@@ Appendix C Baselines
  • SDPO [25] introduces a multi-negative, softmax DPO to better align LLMs with human preferences and mitigate biases in recommendation tasks.
  • Llara [1] proposes aligning large language models with sequential recommenders through hybrid prompt design and curriculum learning for better sequential behavior understanding.
  • SPRec [39] employs a self-play mechanism to iteratively debias preference alignment in LLM-based recommendations by suppressing undesirable items and reinforcing positive samples.
@@
Appendix D Implementation Details
@@
For every LLM variant – including the contrastive-learning baseline (CL, Section 4.3) – we finetune Gemma2-2B-Instruct and Qwen2.5-3B-Instruct. Models are adapted with LoRA (rank = 4) using DeepSpeed Stage 2 and the quantized PagedAdamW optimizer. Training lasts for at most three epochs with early stopping (patience = 1); the maximum generation length is 512 tokens. If public implementations exist, we keep the original hyper-parameters; otherwise, a grid search over learning rates $\{5\times 10^{-4},\,1\times 10^{-4}\}$ is performed.
$\text{R}^{2}\text{ec}$ (ours).
@@
Appendix E Overall Pipeline
@@
7:  Sample a training batch $\mathcal{B}=\{(u,v^{+})\}\sim\mathcal{D}$
@@
Appendix F Gradient Explanation for End-to-End Optimization
We explain how $\text{R}^{2}\text{ec}$ achieves end-to-end optimization by analyzing gradient flow during training, where the recommendation signal propagates through both the reasoning trajectory and the live item encoding process.
During training, each item $v$ is encoded on the fly by the model as
$h_{v}=f_{\theta}(x_{v}),$
so that both the reasoning trajectory producing $h_{T}$ and the item embeddings $h_{v}$ depend on the same evolving computation rather than on static class weights.
At the recommendation step, the model predicts the ground-truth item $v^{+}$ via
$p_{\theta}(v^{+}\mid x_{u},o_{1:T})=\frac{\exp(h_{T}^{\top}h_{v^{+}}/\tau)}{\sum_{v^{\prime}\in B}\exp(h_{T}^{\top}h_{v^{\prime}}/\tau)},$
and optimizes the policy objective for the highest-advantage trajectory $i^{\star}$:
$\mathcal{L}_{\text{rec}}(\theta)=-A_{i^{\star}}\,\log p_{\theta}(v^{+}\mid x_{u},o^{(i^{\star})}_{1:T})$.
The corresponding gradient is
$\nabla_{\theta}\log p_{\theta}(v^{+})=\sum_{v^{\prime}\in B}(\mathbf{1}[v^{\prime}=v^{+}]-p_{\theta}(v^{\prime}))\,\nabla_{\theta}s(v^{\prime}),\quad s(v^{\prime})=h_{T}^{\top}h_{v^{\prime}},$
which expands to
$\nabla_{\theta}s(v^{\prime})=(\nabla_{\theta}h_{T})^{\top}h_{v^{\prime}}+h_{T}^{\top}\nabla_{\theta}f_{\theta}(x_{v^{\prime}}).$
The second component, $h_{T}^{\top}\nabla_{\theta}f_{\theta}(x_{v^{\prime}})$, shows that gradients flow directly into the item encoding process. Hence, updates from the recommendation signal not only refine the reasoning trajectory that produces $h_{T}$, but also reshape the item embedding function $f_{\theta}$, aligning the semantic encoding of items with the reasoning space.
Appendix G Prompt
User Prompt Analyze in depth and finally recommend next {category} I might purchase inside <answer> and </answer> . For example, <answer> a product </answer> . Below is my historical {category} purchases and ratings (out of 5): {% for hist in purchase_histories %} {% {hist.time_delta} ago: [{hist.item_title}] ({hist.rating}) %}
Item Prompt Summarize key attributes of the following {category} inside <answer> and </answer> : {% for key, attr in item.meta %} {% {key}: {attr} %}
@@ Appendix H Limitations
Appendix I Extended Analysis
Scaling to Larger Backbones.
We further scale R2ec to Gemma-9B-it. The results in Table 5 confirm that reasoning consistently improves recommendation across model sizes, and the absolute gain grows with scale (21.7% higher N@5 than its non-reasoning counterpart).
Table 5: Scaling analysis on Gemma backbones. Reasoning consistently improves recommendation performance across model sizes.
Model N@5 N@10 H@5 H@10
Gemma-9B-it
w/o reasoning 0.0370 0.0460 0.0595 0.0893
R2ec 0.0451 0.0527 0.0640 0.0878
Gemma-2B-it
w/o reasoning 0.0321 0.0393 0.0469 0.0692
R2ec 0.0388 0.0457 0.0588 0.0800
Figure 6: Performance comparison across the Electronics (6(a)), GoodReads (6(b)), and MovieLens (6(c)) datasets. GoodReads and MovieLens evaluations follow the setting in SPRec [39] and SDPO [25], respectively, where baseline values are taken directly from their respective original papers.
Method N@5 H@5
GRU4Rec 0.0116 0.0157
Caser 0.0103 0.0164
SASRec 0.0113 0.0167
BigRec 0.0099 0.0146
D3 0.0137 0.0214
LLaRA 0.0183 0.0279
R2ec (ours) 0.0205 0.0314
(a) Electronic
Method N@5 H@5
SASRec 0.0139 0.0202
BigRec 0.0236 0.0310
RW 0.0281 0.0380
D3 0.0324 0.0413
SDPO 0.0315 0.0420
DMPO 0.0314 0.0410
SPRec 0.0330 0.0250
RosePO 0.0215 0.0300
R2ec (ours) 0.0354 0.0561
(b) GoodReads
Method H@1
GRU4Rec 0.3750
Caser 0.3860
SASRec 0.3440
TALLRec 0.3895
MoRec 0.2822
ChatRec 0.2000
LLaRA 0.4730
SDPO 0.5260
R2ec (ours) 0.6211
(c) MovieLens
Cross-Domain Robustness.
To further assess generalization to new domains, we extended our evaluation to the Electronics, MovieLens, and GoodReads datasets, and compared R2ec with advanced baselines, as shown in Tables 6(a), 6(c), and 6(b), respectively. R2ec consistently surpasses prior methods, demonstrating the robustness of reasoning-enhanced recommendation across diverse domains and evaluation settings.
Appendix J Analysis on Case Study
To illustrate domain-specific and backbone-specific reasoning behaviors, we analyze representative cases from the CDs and Video Games datasets (Figures 7, 8, 9, 10) using Gemma and Qwen as backbones. 1) On the CDs dataset, both models tend to reason by discussing several potentially interesting items and explicitly enumerating their features and genres, before providing a final recommendation. This behavior highlights the models’ ability to leverage domain knowledge acquired during pretraining and reinforced through RecPO, for better recommendation. 2) On the Video Games dataset, the models shift towards inferring user intent from the purchase sequence, speculating on the likely category or intended use of the next purchase. This demonstrates a shift from feature-based enumeration to intent-driven reasoning as the domain changes. 3) Notably, Gemma generally produces shorter and more concise reasoning sequences compared to Qwen, which often generates more elaborate justifications. These patterns indicate that, while both models benefit from explicit reasoning, their generation style and reliance on pretrained knowledge can manifest differently across domains and backbone choices.